Circular No. 7404

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 2000bl, 2000bm, 2000bn, 2000bo, 2000bp, 2000bq, 2000br

L.-G. Strolger, University of Michigan and Cerro Tololo Interamerican Observatory (CTIO); R. C. Smith and P. Candia, CTIO; A. Soderberg, Bates College; P. Challis, Harvard-Smithsonian Center for Astrophysics; and the rest of the Nearby Galaxies Supernova Search Team report the discovery of seven new supernovae in their recent search campaign conducted Mar. 21–31 using the Kitt Peak 0.9-m telescope (+ Mosaic camera). The search is conducted to limiting mag R ∼ 21 in ∼ 250 1^{o}-square fields along the celestial equator. Search epoch frames are compared to template frames obtained from two weeks to a year previously.

SN  2000 UT  α_{2000}  δ_{2000}  R  Offset
2000bl Mar. 24 8^{h}55^{m}20.^{s}44 − 7^{o}23′42′′.9 19.4 0′′.3 E, 1′′.3 N
2000bm Mar. 23 8 58 56.02 + 5 10 15.7 19.2 0′′.6 W, 2′′.1 N
2000bn Mar. 25 9 35 41.15 + 4 32 13.3 18.8  —
2000bo Mar. 23 9 49 34.92 − 3 20 38.3 18.3 2′′.1 E, 1′′.3 S
2000bp Mar. 25 9 54 20.63 − 4 51 13.3 19.8 0′′.4 E, 0′′.3 N
2000bq Mar. 25 11 01 03.24 − 5 50 36.2 20.4  —
2000br Mar. 25 14 03 50.20 − 4 01 01.7 19.7 0′′.3 E, 5′′.5 N

R. Corvarrubias (CTIO) obtained confirming CCD images of SNe 2000bm, 2000bq, and 2000br on Mar. 29 using the CTIO 0.9-m telescope. J. Krick and A. Athey (University of Michigan) obtained confirming CCD images of SNe 2000bn and 2000bp on Apr. 1, and of SNe 2000bl and 2000bo on Apr. 2, using the Hiltner 2.4-m telescope at MDM Observatory. SN 2000bq has no visible host galaxy. SN 2000bn is at the host galaxy’s center. M. M. Phillips (Las Campanas Observatory) identifies SN 2000bn as a type-Ia supernova, at or before maximum light, from spectra (range 400–800 nm) obtained by Strolger and Smith at the Mayall 4-m telescope (+ R-C Spectrograph) on the nights of Apr. 2–4. From spectra obtained on the same evenings by Strolger and Smith, Phillips identifies SN 2000bp as a young type-II supernova with a redshift of ∼ 0.06. SN 2000br is identified as a type-II supernova from spectra obtained at the Mayall 4-m telescope by R. Knopp and A. Conely (Lawrence Berkeley Laboratories) on Apr. 1.83 UT and by Strolger and Smith on Apr. 2–4.

2000 April 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green